# Probing the accretion geometry of black hole X-ray binaries: A multi-mission spectro-polarimetric and timing study

### v1

@@ 5.4 Frontiers of X-ray Polarimetry: A Promising Era Ahead
Figure 8: (a) Time binned ($50$ s) light curve of Swift J1727.816131727.816131727.8-16131727.8 - 1613 observed with IXPE in the hard-intermediate state (HIMS). (b) Best-fitted power density spectra ($0.1-50$ Hz) obtained from the quasi-simultaneous NICER ($0.5-10$ keV), NuSTAR ($3-20$ keV) and AstroSat ($20-100$ keV) observations. (c) Best-fitted energy spectra from the quasi-simultaneous NICER and NuSTAR observations in $0.7-40$ keV energy range. (d) IXPE image ($2-8$ keV) of Swift J1727.816131727.816131727.8-16131727.8 - 1613. The red circular boundary denotes $60$ arcsec source region considered for polarimetric analysis. (e) Variation of the polarization degree (PD) with energy obtained from the model-independent polarimetric analysis. The gray histograms represent the $\rm MDP_{\rm 99}\%$ level. See the text for details.
@@
Indeed, the first X-ray polarization measurements were conducted in the 1970s by OSO-8, revealing significant polarization in Cyg X11-1- 1 with PD $\sim 2.4\%$ at $2.6$ keV and $\sim 5.3\%$ at $5.2$ keV (Long et al., 1980). Thereafter, higher polarization was observed in hard X-rays for Cyg X11-1- 1 with INTEGRAL as $\lesssim 20\%$ in the $130-230$ keV band (Jourdain et al., 2012) and $\sim 75\%$ above $400$ keV (Rodriguez et al., 2015). Recently, Chattopadhyay et al. (2024) detected $\sim 23\%$ polarization in the HIMS of Cyg X11-1- 1 using AstroSat/CZTI. Despite these advances, low-energy X-ray polarimetry remained relatively underexplored until the launch of IXPE, which has made substantial progress in this domain during its first three and a half years of operation. Nevertheless, several fundamental questions remain unanswered, paving the way for future research.
@@
At present, Swift J1727.816131727.816131727.8-16131727.8 - 1613 is the only BH-XRB transient observed during its outburst with IXPE exhibiting polarization over four spectral states (HIMS, SIMS, HSS and LHS) along with QPO features. To establish robust spectro-polarimetric correlations across different spectral states, coordinated IXPE observations along with simultaneous spectro-temporal coverage from other missions are indispensable. Further, since the effects of Comptonization become prominent beyond the coverage of IXPE up to $8$ keV, complementary wide-band X-ray spectro-polarimetry and timing studies are crucial to constrain the coronal geometry. Thus, the four windows (see Fig. 8) of X-ray astronomy (imaging, timing, spectroscopy and polarimetry) with wide energy coverage is crucial to unravel the complete understanding of BH-XRBs (see Fig. 8).
@@
In this context, the recently launched XPoSat [^16] https://www.isro.gov.in/XPoSat.html mission and the upcoming XL-Calibur mission (Abarr et al., 2021) are set to play a crucial role in investigating hard X-ray polarimetric properties of BH-XRBs. In particular, POLIX ($8-30$ keV) and XSPECT ($0.8-15$ keV) onboard on XPoSat are suitable for mid-energy polarimetry and simultaneous spectro-temporal studies, respectively. The balloon-borne XL-Calibur mission, operating in $15-80$ keV energy range, completed a successful flight [^17] https://sscspace.com/nasa-xl-calibur-balloon-launched/ in July 2024, aiming to measure polarization in Cyg X11-1- 1. Additionally, the future soft X-ray spectro-polarimetric mission eXTP [^18] https://heasarc.gsfc.nasa.gov/docs/heasarc/missions/extp.html ($2-10$ keV), expected to launch in $2027$, will offer improved sensitivity for low-energy polarization measurements. Needless to mention that these dedicated missions will significantly advance our understanding of the accretion dynamics and coronal geometries in BH-XRBs. However, simultaneous observations combining wide-band timing and spectro-polarimetric measurement are still lacking, which emphasizes the need for future missions with broader energy ranges and enhanced capabilities.
@@ 6 Conclusions
  • The combined spectro-temporal results in the wide-band energy range ($0.5-100$ keV) reveal the presence of distinct canonical spectral states of BH-XRBs, which are closely connected to the temporal characteristics and emission mechanism of the sources. The timing and spectral features result in degeneracy among the different viable disc-corona geometries of the sources.
@@
  • Timing and spectro-polarimetric results, combined with known radio-jet angles, suggest that a two-component disc-corona model may be plausible. In this framework, a radially extended strong corona is likely to be present at the inner part of a truncated accretion disc during harder states, while a comparatively weaker corona persists in softer states for Swift J1727.8--1613, Cyg X--1, GX 339--4, and IGR J17091--3624.
  • For 4U 1630--47, LMC X--3, and 4U 1957$+$115, a thermally dominated accretion disc with a weak corona seems to be preferred. However, the coronal geometry remains unclear due to the minimal contribution of Comptonized emission in the IXPE band and the lack of complementary radio observations, though this analysis does support the presence of disc winds in 4U 1630--47. We also note that the disc-corona geometry outlined in this work bears limitations to explain the observed features of Cyg X--3, for which an alternative configuration involving an optically thick inner funnel has been proposed (Veledina et al., 2024a, b).
@@ 7 Acknowledgments
SM and SD thank the Department of Physics, IIT Guwahati, India for providing the facilities to complete this work. SM also thanks Luca Baldini for the discussion in identifying the bugs in IXPEOBSSIM software and releasing the latest version IXPEOBSSIM V31.0.3. AK, SS, KJ and AN thank GH, SAG; DD, PDMSA, and Director, URSC for encouragement and continuous support to carry out this research. This publication uses data from the IXPE, NICER, NuSTAR and AstroSat missions. This research has made use of the MAXI data provided by RIKEN, JAXA and the MAXI team (Matsuoka et al., 2009). The Swift/BAT transient monitor results provided by the Swift/BAT team are also used (Krimm et al., 2013). We thank each instrument team for processing the data and providing the necessary software tools for the analysis.
@@ 8 Data Availability
($2-20$ keV) ($15-50$ keV)
Cyg X11-1- 1 E1 IXPE 01002901 2022-05-15 59714.64 242 $9.82\pm 0.21$ $401\pm 11$ $928\pm 34$- LHS
NICER 5100320101 2022-05-15 59714.26 8.2 $2938\pm 25$- -- -- --
NuSTAR 30702017002 2022-05-18 59717.57 16 $324\pm 15$- -- -- --
AstroSat A11_080T01_9000005146 2022-05-15 59714.03 $47.3$ $1505\pm 37$- -- -- --
@@
4U 16304747-47- 47 E1 IXPE 01250401 2022-08-23 59814.95 459 $4.56\pm 0.08$ $247\pm 8$ $10\pm 2$- HSS
NICER 5501010102 2022-08-23 59814.01 2 $479\pm 10$- -- -- --
NuSTAR 80802313002 2022-08-25 59816.19 17 $348\pm 14$- -- -- --
@@
NICER 6557010101 2023-03-10 60013.76 5 $1009\pm 15$- -- -- --
NuSTAR 80801327002 2023-03-09 60012.36 12 $884\pm 29$- -- -- --
AstroSat A12_056T02_9000005538 2023-03-10 60013.08 29.4 $1938\pm 43$- -- -- --
Cyg X33-3- 3 E1 IXPE 02001899 2022-10-14 59866.04 538 $1.33\pm 0.04$ $106\pm 20$ $129\pm 26$ $106\pm 24^{*}$ LHS
NICER 5142010105 2022-10-14 59866.28 5 $44.79\pm 3.12$- -- -- --
NuSTAR 90802323002 2022-10-13 59865.62 18 $197\pm 1$- -- -- --
@@
LMC X11-1- 1 E1 IXPE 02001901 2022-10-19 59871.61 563 $0.96\pm 0.03$ $14\pm 9$ $13\pm 3$- HSS
NICER 5100070101 2022-10-19 59871.79 2.8 $193\pm 6$- -- -- --
NuSTAR 90801324002 2022-10-24 59876.16 19 $15\pm 2$- -- -- --
4U 1957$+115$ E1 IXPE 02006601 2023-05-12 60076.10 572 $1.09\pm 0.04$ $19\pm 4$ $2\pm 1$- HSS
NICER 6100400101 2023-05-12 60076.55 0.6 $265\pm 8$- -- -- --
NuSTAR 30902042002 2023-05-15 60079.33 19 $29\pm 4$- -- -- --
LMC X33-3- 3 E1 IXPE 02006599 2023-07-07 60132.77 562 $0.97\pm 0.03$ $16\pm 4$ $18\pm 3$- HSS
NICER 6101010117 2023-07-08 60133.42 0.1 $358\pm 23$- -- -- --
NuSTAR 30902041002 2023-07-09 60134.51 28 $15\pm 3$- -- -- --
@@
NICER 7704010101 2024-10-03 60586.63 2.4 $606\pm 11$- -- -- --
E3 IXPE 03004901 2024-11-21 60635.94 382 $0.42\pm 0.02$ $8\pm 5$- -- HSS
NICER 7704010201 2024-11-23 60637.10 1.7 $246\pm 7$- -- -- --
@@
($2-20$ keV) ($15-50$ keV)
Swift J1727.816131613-1613- 1613 E1 IXPE 02250901 2023-09-07 60194.81 19 $42.21\pm 0.28$ $6053\pm 71$ $4898\pm 173$ $120\pm 12^{**}$ HIMS
NICER 6203980115 2023-09-08 60195.61 0.2 $64955\pm 274$- -- -- --
NuSTAR 80902333006 2023-09-07 60194.78 0.7 $6536\pm 8$- -- -- --
AstroSat T05_145T01_9000005836 2023-09-08 60195.07 $32$ $541\pm 23$- -- -- --
@@
NICER 6203980119 2023-09-13 60200.18 4.9 $65392\pm 122$- -- -- --
NuSTAR 80902313002 2023-09-16 60203.81 0.6 $5478\pm 26$- -- -- --
E3 IXPE 02251101 2023-09-27 60214.92 21 $36.85\pm 0.22$ $3395\pm 49$ $300\pm 10$ $88\pm 2.64^{*}$ HIMS
@@
NICER 7708010109 2024-04-03 60403.29 2.5 $675\pm 12$- -- -- --
GX $339-4$ E1 IXPE 03005101 2024-02-14 60354.93 95 $17.44\pm 0.21$- $79\pm 25$- SIMS
NICER 7702010112 2024-02-14 60354.98 3.8 $5263\pm 34$- -- -- --
NuSTAR 91002306002 2024-02-14 60354.70 16 $308\pm 13$- -- -- --
AstroSat A05_166T01_9000006070 2024-02-14 60354.31 $50.7$ $629\pm 26$- -- -- --
@@
AstroSat A13_028T01_9000006122 2024-03-10 60379.44 -- -- -- -- -- --
Swift J151857.0572147151857.0572147151857.0-572147151857.0 - 572147 E1 IXPE 03250201 2024-03-18 60387.15 96 $12.18\pm 0.12$ $422\pm 8$ $152\pm 17$- HSS
NICER 7204220111 2024-03-18 60387.55 4.2 $1468\pm 18$- -- -- --
NuSTAR 91001311004 2024-03-18 60387.64 9.2 $444\pm 17$- -- -- --
IGR J17091362417091362417091-362417091 - 3624 E1 IXPE 04250201 2025-03-07 60741.30 163 $0.27\pm 0.02$ $755\pm 11$ $82\pm 27$- LHS
NuSTAR 81002342008 2025-03-07 60741.58 21 $27\pm 15$- -- -- --
MAXI J174429417442941744-2941744 - 294 E1 IXPE 04250301 2025-04-05 60770.09 149 $0.17\pm 0.11$ $45\pm 28$- -- HSS
@@
LMC X33-3- 3 E1 $0.04^{*}$ $1.10_{-0.01}^{+0.01}$ $10^{*}$ $1.78_{-0.35}^{+0.45}$ $0.005_{-0.002}^{+0.003}$ $567/543$ $0.061$ $0.001$ $0.062$ $13.4$ HSS $0.7-20$-
($d=48.1$ kpc) E2 $0.04^{*}$ $1.09_{-0.01}^{+0.01}$ $10^{*}$ $2.51_{-0.07}^{+0.08}$ $0.3^{*}$ $153/128$ $0.11$ $0.03$ $0.14$ $30.8$ HSS $1-10$-
E3 $0.04^{*}$ $0.6_{-0.03}^{+0.04}$ $1.04_{-0.05}^{+0.08}$ $1.72_{-0.08}^{+0.11}$ $0.4^{*}$ $118/117$ $0.03$ $0.006$ $0.035$ $7.7$ HSS $1-10$-
Swift J1727.816131613-1613- 1613 E1 $0.23_{-0.01}^{+0.01}$ $0.29_{-0.02}^{+0.01}$ $8.44_{-0.11}^{+0.16}$ $1.98_{-0.02}^{+0.04}$ $0.74_{-0.01}^{+0.01}$ $1181/931$ $10.07$ $15.01$ $25.24$ $17.5$ HIMS $0.7-40$ $12$
($d=2.7$ kpc) E2 $0.20_{-0.01}^{+0.01}$ $0.43_{-0.02}^{+0.02}$ $9.34_{-0.23}^{+0.22}$ $2.11_{-0.04}^{+0.05}$ $0.82_{-0.02}^{+0.02}$ $912/847$ $11.35$ $13.58$ $25.11$ $17.4$ HIMS $0.7-40$-
E4 $0.25_{-0.01}^{+0.01}$ $0.85_{-0.02}^{+0.03}$ $10^{*}$ $2.74_{-0.04}^{+0.04}$ $0.38_{-0.03}^{+0.04}$ $1373/1447$ $15.89$ $3.63$ $25.04$ $17.3$ SIMS $0.7-40$ $6.1$
E5 $0.29_{-0.01}^{+0.02}$ $0.88_{-0.02}^{+0.02}$ $10^{*}$ $2.32_{-0.04}^{+0.06}$ $0.19_{-0.02}^{+0.02}$ $1572/1601$ $10.12$ $2.09$ $16.78$ $11.6$ SIMS $0.7-60$ $6.4$
E6 $0.17_{-0.01}^{+0.01}$ $0.48_{-0.01}^{+0.01}$ $10^{*}$ $5.14_{-0.39}^{+0.36}$ $0.17_{-0.05}^{+0.03}$ $127/132$ $0.91$ $0.02$ $0.93$ $0.6$ HSS $0.7-10$ $1.4$
E7 $0.18_{-0.01}^{+0.01}$ $0.45_{-0.01}^{+0.01}$ $10^{*}$ $4.57_{-0.60}^{+0.65}$ $0.13_{-0.05}^{+0.09}$ $102/111$ $0.66$ $0.01$ $0.67$ $0.5$ HSS $0.7-10$ $1.3$
E8 $0.12^{*}$ $0.24_{-0.01}^{+0.01}$ $20^{*}$ $1.71_{-0.02}^{+0.03}$ $0.35_{-0.01}^{+0.01}$ $126/130$ $0.08$ $0.12$ $0.19$ $0.1$ LHS $0.7-10$ $30.2$
GX $339-4$ E1 $0.51_{-0.01}^{+0.01}$ $0.60_{-0.02}^{+0.02}$ $0.97_{-0.01}^{+0.02}$ $2.02_{-0.52}^{+0.44}$ $0.34_{-0.05}^{+0.08}$ $1603/1538$ $0.84$ $0.36$ $1.71$ $11.5$ SIMS $0.7-50$ $14.2$
@@
  • $L_{\rm Edd}=1.26\times 10^{39}$ erg $\rm s^{-1}$ for $10M_{\odot}$ BH. Calculated in $1-100$ keV energy range. Computed in $0.1-50$ Hz using NICER observations in $0.5-10$ keV energy range. Frozen values. Results are obtained using model M3.
@@
bGX $339-4$ E1 $0.85^{*}$ $50^{*}$ $2.27_{-0.03}^{+0.03}$ $0.5^{\boxtimes}$ $3.92_{-0.05}^{+0.12}$ $3.16_{-0.33}^{+0.35}$ $0.54$ SIMS
cSwift J1727.816131727.816131727.8-16131727.8 - 1613 E4 $0.98^{*}$ $40^{*}$ $2.19_{-0.02}^{+0.03}$ $5.02^{\boxtimes}$ $4.11_{-0.13}^{+0.12}$ $0.68_{-0.14}^{+0.17}$ $5.44$ SIMS

### v2

@@ 5.4 Frontiers of X-ray Polarimetry: A Promising Era Ahead
Figure 8: (a) Time binned ($50$ s) light curve of Swift J1727.816131727.8-1613 observed with IXPE in the hard-intermediate state (HIMS). (b) Best-fitted power density spectra (0.1500.1-50 Hz) obtained from the quasi-simultaneous NICER (0.5100.5-10 keV), NuSTAR (3203-20 keV) and AstroSat (2010020-100 keV) observations. (c) Best-fitted energy spectra from the quasi-simultaneous NICER and NuSTAR observations in 0.7400.7-40 keV energy range. (d) Colour composite X-ray image (NuSTAR: reddish and IXPE: greenish) of Swift J1727.816131727.8-1613 with the overlap in yellow colour. The magenta cross marks the source location in X-ray and the cyan circle of radius 60 arcsec around it denotes the region considered for polarimetric analysis. (e) Variation of the polarization degree (PD) with energy obtained from the model-independent polarimetric analysis. The gray histograms represent the $\rm MDP_{\rm 99}\%$ level.
@@
Indeed, the first X-ray polarization measurements were conducted in the 1970s by OSO-8, revealing significant polarization in Cyg X1-1 with PD $\sim 2.4\%$ at $2.6$ keV and $\sim 5.3\%$ at $5.2$ keV (Long et al., 1980). Thereafter, higher polarization was observed in hard X-rays for Cyg X1-1 with INTEGRAL as $\lesssim 20\%$ in the 130230130-230 keV band (Jourdain et al., 2012) and $\sim 75\%$ above $400$ keV (Rodriguez et al., 2015). Recently, Chattopadhyay et al. (2024) detected $\sim 23\%$ polarization in the HIMS of Cyg X1-1 using AstroSat/CZTI. Despite these advances, low-energy X-ray polarimetry remained relatively underexplored until the launch of IXPE, which has made substantial progress in this domain during its first three and a half years of operation. Nevertheless, several questions remain unanswered, paving the way for future research.
@@
At present, Swift J1727.816131727.8-1613 is the only BH-XRB transient observed during its outburst with IXPE exhibiting polarization over four spectral states (HIMS, SIMS, HSS and LHS) along with QPO features. To establish robust spectro-polarimetric correlations across different spectral states, coordinated IXPE observations along with simultaneous spectro-temporal coverage from other missions are indispensable. Further, since the effects of Comptonization become prominent beyond the coverage of IXPE up to $8$ keV, complementary wide-band X-ray spectro-polarimetry and timing studies are crucial to constrain the coronal geometry. Thus, the four windows of X-ray astronomy (imaging, timing, spectroscopy and polarimetry) with wide energy coverage is crucial to unravel the complete understanding of BH-XRBs (see Fig. 8).
@@
In this context, the recently launched XPoSat [^10] https://www.isro.gov.in/XPoSat.html mission and the upcoming XL-Calibur mission (Abarr et al., 2021) are set to play a crucial role in investigating hard X-ray polarimetric properties of BH-XRBs. In particular, POLIX (8308-30 keV) and XSPECT (0.8150.8-15 keV) onboard on XPoSat are suitable for mid-energy polarimetry and simultaneous spectro-temporal studies, respectively. The balloon-borne XL-Calibur mission, operating in 158015-80 keV energy range, completed a successful flight [^11] https://sscspace.com/nasa-xl-calibur-balloon-launched/ in July 2024, aiming to measure polarization in Cyg X1-1. This eventually led to the recent measurement of a polarization degree of $\sim 5\%$ in LHS of Cyg X1-1 in $\sim 19-64$ keV energy range (Awaki et al., 2025), which is comparable to the low energy (282-8 keV) polarimetric results of Cyg X1-1 (PD $\sim 4.8\%$, LHS) with IXPE (see Fig. 4 and Table 5). The consistency of the measured PDs and PAs across the 2642-64 keV range observed with IXPE and XL-Calibur indicates that the same underlying physical processes likely govern the observed polarization in both energy bands (Awaki et al., 2025). In addition, the future soft X-ray spectro-polarimetric mission eXTP [^12] https://heasarc.gsfc.nasa.gov/docs/heasarc/missions/extp.html (2102-10 keV), expected to launch in $2027$, will offer improved sensitivity for low-energy polarization measurements. Indeed, these dedicated missions will significantly advance our understanding of the accretion dynamics and coronal geometries in BH-XRBs. However, simultaneous observations combining wide-band timing and spectro-polarimetric measurement are still lacking, which emphasizes the need for future missions with broader energy ranges and enhanced capabilities.
@@ 6 Conclusions
  • The combined spectro-temporal results in the wide-band energy range (0.51000.5-100 keV) reveal the presence of distinct canonical spectral states of BH-XRBs, which are closely connected to the temporal characteristics and emission mechanism of the sources. The timing and spectral features result in degeneracy among the different viable disc-corona geometries of the sources.
@@
  • Timing and spectro-polarimetric results, combined with known radio-jet angles, suggest that a two-component disc-corona model may be plausible. In this framework, a radially extended strong corona is likely to be present at the inner part of a truncated accretion disc during harder states, while a comparatively weaker corona persists in softer states for Swift J1727.8-1613, Cyg X-1, GX 339-4, and IGR J17091-3624.
  • For 4U 1630-47, LMC X-3, and 4U 1957++115, a thermally dominated accretion disc with a weak corona seems to be preferred. However, the coronal geometry remains unclear due to the minimal contribution of Comptonized emission in the IXPE band and the lack of complementary radio observations, though this analysis does support the presence of disc winds in 4U 1630-47. We also note that the disc-corona geometry outlined in this work bears limitations to explain the observed features of Cyg X-3, for which an alternative configuration involving an optically thick inner funnel has been proposed (Veledina et al., 2024a, b).
@@ 7 Acknowledgments
Authors thank the anonymous reviewer for constructive comments and useful suggestions that helped to improve the quality of the manuscript. SM and SD thank the Department of Physics, IIT Guwahati, India for providing the facilities to complete this work. SM also thanks Luca Baldini for the discussion in identifying the bugs in IXPEOBSSIM software and releasing the latest version IXPEOBSSIM V31.0.3. AK, SS, KJ and AN thank GH, SAG; DD, PDMSA, and Director, URSC for encouragement and continuous support to carry out this research. This publication uses data from the IXPE, NICER, NuSTAR and AstroSat missions. This research has made use of the MAXI data provided by RIKEN, JAXA and the MAXI team (Matsuoka et al., 2009). The Swift/BAT transient monitor results provided by the Swift/BAT team are also used (Krimm et al., 2013). We thank each instrument team for processing the data and providing the necessary software tools for the analysis.
@@ 8 Data Availability
(2202-20 keV) (155015-50 keV)
Cyg X1-1 E1 IXPE 01002901 2022-05-15 59714.64 242 $9.82\pm 0.21$ $401\pm 11$ $928\pm 34$ - LHS
NICER 5100320101 2022-05-15 59714.26 8.2 $2938\pm 25$ - - - -
NuSTAR 30702017002 2022-05-18 59717.57 16 $324\pm 15$ - - - -
AstroSat A11_080T01_9000005146 2022-05-15 59714.03 $47.3$ $1505\pm 37$ - - - -
@@
4U 163047-47 E1 IXPE 01250401 2022-08-23 59814.95 459 $4.56\pm 0.08$ $247\pm 8$ $10\pm 2$ - HSS
NICER 5501010102 2022-08-23 59814.01 2 $479\pm 10$ - - - -
NuSTAR 80802313002 2022-08-25 59816.19 17 $348\pm 14$ - - - -
@@
NICER 6557010101 2023-03-10 60013.76 5 $1009\pm 15$ - - - -
NuSTAR 80801327002 2023-03-09 60012.36 12 $884\pm 29$ - - - -
AstroSat A12_056T02_9000005538 2023-03-10 60013.08 29.4 $1938\pm 43$ - - - -
Cyg X3-3 E1 IXPE 02001899 2022-10-14 59866.04 538 $1.33\pm 0.04$ $106\pm 20$ $129\pm 26$ $106\pm 24^{*}$ LHS
NICER 5142010105 2022-10-14 59866.28 5 $44.79\pm 3.12$ - - - -
NuSTAR 90802323002 2022-10-13 59865.62 18 $197\pm 1$ - - - -
@@
LMC X1-1 E1 IXPE 02001901 2022-10-19 59871.61 563 $0.96\pm 0.03$ $14\pm 9$ $13\pm 3$ - HSS
NICER 5100070101 2022-10-19 59871.79 2.8 $193\pm 6$ - - - -
NuSTAR 90801324002 2022-10-24 59876.16 19 $15\pm 2$ - - - -
4U 1957+115+115 E1 IXPE 02006601 2023-05-12 60076.10 572 $1.09\pm 0.04$ $19\pm 4$ $2\pm 1$ - HSS
NICER 6100400101 2023-05-12 60076.55 0.6 $265\pm 8$ - - - -
NuSTAR 30902042002 2023-05-15 60079.33 19 $29\pm 4$ - - - -
LMC X3-3 E1 IXPE 02006599 2023-07-07 60132.77 562 $0.97\pm 0.03$ $16\pm 4$ $18\pm 3$ - HSS
NICER 6101010117 2023-07-08 60133.42 0.1 $358\pm 23$ - - - -
NuSTAR 30902041002 2023-07-09 60134.51 28 $15\pm 3$ - - - -
@@
NICER 7704010101 2024-10-03 60586.63 2.4 $606\pm 11$ - - - -
E3 IXPE 03004901 2024-11-21 60635.94 382 $0.42\pm 0.02$ $8\pm 5$ - - HSS
NICER 7704010201 2024-11-23 60637.10 1.7 $246\pm 7$ - - - -
E4 IXPE 03005001 2025-01-11 60686.18 389 $0.86\pm 0.41$ $11\pm 4$ $9\pm 7$ - HSS
NuSTAR 31002032006 2025-01-11 60686.68 35 $16\pm 5$ - - -
@@
(2202-20 keV) (155015-50 keV)
Swift J1727.81613-1613 E1 IXPE 02250901 2023-09-07 60194.81 19 $42.21\pm 0.28$ $6053\pm 71$ $4898\pm 173$ $120\pm 12^{**}$ HIMS
NICER 6203980115 2023-09-08 60195.61 0.2 $64955\pm 274$ - - - -
NuSTAR 80902333006 2023-09-07 60194.78 0.7 $6536\pm 8$ - - - -
AstroSat T05_145T01_9000005836 2023-09-08 60195.07 $32$ $541\pm 23$ - - - -
@@
NICER 6203980119 2023-09-13 60200.18 4.9 $65392\pm 122$ - - - -
NuSTAR 80902313002 2023-09-16 60203.81 0.6 $5478\pm 26$ - - - -
E3 IXPE 02251101 2023-09-27 60214.92 21 $6.85\pm 0.22 33$$395\pm 49$ $300\pm 10$ $88\pm 2.64^{*}$ HIMS
@@
NICER 7708010109 2024-04-03 60403.29 2.5 $675\pm 12$ - - - -
GX 3394339-4 E1 IXPE 03005101 2024-02-14 60354.93 95 $17.44\pm 0.21$ - $79\pm 25$ - SIMS
NICER 7702010112 2024-02-14 60354.98 3.8 $5263\pm 34$ - - - -
NuSTAR 91002306002 2024-02-14 60354.70 16 $308\pm 13$ - - - -
AstroSat A05_166T01_9000006070 2024-02-14 60354.31 $50.7$ $629\pm 26$ - - - -
@@
AstroSat A13_028T01_9000006122 2024-03-10 60379.44 - - - - - -
Swift J151857.0572147151857.0-572147 E1 IXPE 03250201 2024-03-18 60387.15 96 $12.18\pm 0.12$ $422\pm 8$ $152\pm 17$ - HSS
NICER 7204220111 2024-03-18 60387.55 4.2 $1468\pm 18$ - - - -
NuSTAR 91001311004 2024-03-18 60387.64 9.2 $444\pm 17$ - - - -
IGR J17091362417091-3624 E1 IXPE 04250201 2025-03-07 60741.30 163 $0.27\pm 0.02$ $755\pm 11$ $82\pm 27$ - LHS
NuSTAR 81002342008 2025-03-07 60741.58 21 $27\pm 15$ - - - -
MAXI J17442941744-294 E1 IXPE 04250301 2025-04-05 60770.09 149 $0.17\pm 0.11$ $45\pm 28$ - - HSS
GRS $1915+105$ E1 IXPE 04003501 2025-05-10 60805.69 142 - $6\pm 3$ - - -
@@
LMC X3-3 E1 $0.04^{*}$ $1.10_{-0.01}^{+0.01}$ $10^{*}$ $1.78_{-0.35}^{+0.45}$ $0.005_{-0.002}^{+0.003}$ $567/543$ $0.061$ $0.001$ $0.062$ $13.4$ HSS 0.7200.7-20 -
($d=48.1$ kpc) E2 $0.04^{*}$ $9_{-0.01}^{+0.01} 1010^{*}$ $2.51_{-0.07}^{+0.08}$ $0.3^{*}$ $153/128$ $0.11$ $0.03$ $0.14$ $30.8$ HSS 1101-10 -
E3 $0.04^{*}$ $0.6_{-0.03}^{+0.04}$ $1.04_{-0.05}^{+0.08}$ $1.72_{-0.08}^{+0.11}$ $0.4^{*}$ $118/117$ $0.03$ $0.006$ $0.035$ $7.7$ HSS 1101-10 -
E4 $0.04^{*}$ $7_{-0.01}^{+0.01} 1010^{*}$ $2.05_{-0.04}^{+0.05}$ $0.005_{-0.003}^{+0.003}$ $771/706$ $0.06$ $0.006$ $0.07$ $15.4$ HSS 3303-30 -
Swift J1727.81613-1613 E1 $0.23_{-0.01}^{+0.01}$ $0.29_{-0.02}^{+0.01}$ $8.44_{-0.11}^{+0.16}$ $1.98_{-0.02}^{+0.04}$ $0.74_{-0.01}^{+0.01}$ $1181/931$ $10.07$ $15.01$ $25.24$ $17.5$ HIMS 0.7400.7-40 $12$
($d=2.7$ kpc) E2 $0.20_{-0.01}^{+0.01}$ $0.43_{-0.02}^{+0.02}$ $9.34_{-0.23}^{+0.22}$ $2.11_{-0.04}^{+0.05}$ $0.82_{-0.02}^{+0.02}$ $912/847$ $11.35$ $13.58$ $25.11$ $17.4$ HIMS 0.7400.7-40 -
E4 $0.25_{-0.01}^{+0.01}$ $0.85_{-0.02}^{+0.03}$ $10^{*}$ $2.74_{-0.04}^{+0.04}$ $0.38_{-0.03}^{+0.04}$ $1373/1447$ $15.89$ $3.63$ $25.04$ $17.3$ SIMS 0.7400.7-40 $6.1$
E5 $0.29_{-0.01}^{+0.02}$ $0.88_{-0.02}^{+0.02}$ $10^{*}$ $2.32_{-0.04}^{+0.06}$ $0.19_{-0.02}^{+0.02}$ $1572/1601$ $10.12$ $2.09$ $16.78$ $11.6$ SIMS 0.7600.7-60 $6.4$
E6 $0.17_{-0.01}^{+0.01}$ $0.48_{-0.01}^{+0.01}$ $10^{*}$ $5.14_{-0.39}^{+0.36}$ $0.17_{-0.05}^{+0.03}$ $127/132$ $0.91$ $0.02$ $0.93$ $0.6$ HSS 0.7100.7-10 $1.4$
E7 $0.18_{-0.01}^{+0.01}$ $0.45_{-0.01}^{+0.01}$ $10^{*}$ $4.57_{-0.60}^{+0.65}$ $0.13_{-0.05}^{+0.09}$ $102/111$ $0.66$ $0.01$ $0.67$ $0.5$ HSS 0.7100.7-10 $1.3$
E8 $0.12^{*}$ $4_{-0.01}^{+0.01} 2020^{*}$ $1.71_{-0.02}^{+0.03}$ $0.35_{-0.01}^{+0.01}$ $126/130$ $0.08$ $0.12$ $0.19$ $0.1$ LHS 0.7100.7-10 $30.2$
GX 3394339-4 E1 $0.51_{-0.01}^{+0.01}$ $0.60_{-0.02}^{+0.02}$ $0.97_{-0.01}^{+0.02}$ $2.02_{-0.52}^{+0.44}$ $0.34_{-0.05}^{+0.08}$ $1603/1538$ $0.84$ $0.36$ $1.71$ $11.5$ SIMS 0.7500.7-50 $14.2$
@@
  • $L_{\rm Edd}=1.26\times 10^{39}$ erg $\rm s^{-1}$ for $10M_{\odot}$ BH. Calculated in 11001-100 keV energy range. Computed in 0.1500.1-50 Hz using NICER observations in 0.5100.5-10 keV energy range. Frozen values. Results are obtained using model M3.
@@
bGX 3394339-4 E1 $0.85^{*}$ $50^{*}$ $2.27_{-0.03}^{+0.03}$ $0.5^{\boxtimes}$ $3.92_{-0.05}^{+0.12}$ $3.16_{-0.33}^{+0.35}$ $0.54$ SIMS
cSwift J1727.816131727.8-1613 E4 $0.98^{*}$ $40^{*}$ $2.19_{-0.02}^{+0.03}$ $5.02^{\boxtimes}$ $4.11_{-0.13}^{+0.12}$ $0.68_{-0.14}^{+0.17}$ $5.44$ SIMS